OpenStudy (anonymous):

A transverse sinusoidal wave on a string has a period T = 35.0 ms and travels in the negative x direction with a speed of 30.0 m/s. At t = 0, a particle on the string at x = 0 has a transverse position of 2.00 cm and is traveling downward with a speed of 5.50 m/s. (a) What is the amplitude of the wave? (b) What is the phase constant? (c) What is the maximum transverse speed of the string? I've tried it all and always seem to get the wrong answer. I'm looking for some guidance.

OpenStudy (anonymous):

Givens: \[T=0.035 s\]\[y(0,0)=0.02m\]\[v=30.0 m/s\]\[v_y=-5.50 m/s\]found:\[f=1/T=28.57Hz\]\[\lambda=v/f=1.05m\]\[\omega=2pi/T=179.52 rads/s\]

OpenStudy (anonymous):

Finding the amplitude of the wave will require me to find

OpenStudy (anonymous):

Phi